OpenStudy (anonymous):

How many seconds Re in 1/6 of a min

OpenStudy (anonymous):

ARE*

OpenStudy (anonymous):

490 sec... 8 1/6=((8*6)+1)/6=49/6 i min has 60 sec... so 49/6 min has=49/6*60=490...